# NTK-DFL: Enhancing Decentralized Federated Learning in Heterogeneous Settings via Neural Tangent Kernel

### v1

@@ 5 Conclusion and Future Work
In this paper, we have introduced NTK-DFL, a novel approach to decentralized federated learning that leverages the neural tangent kernel to address the challenges of statistical heterogeneity in decentralized learning settings. Our method combines the expressiveness of NTK-based weight evolution with a decentralized architecture, allowing for efficient and privacy-preserving collaborative learning without a central server. We reduce the number of communication rounds needed for convergence, which may prove advantageous for high-latency settings or those with heavy encoding/decoding costs.
There are promising unexplored directions for NTK-DFL. For instance, extending the algorithm to training models such as CNNs, ResNets (He et al., 2016), and transformers (Vaswani et al., 2017). Additionally, future research could explore the application of NTK-DFL to cross-silo federated learning scenarios, particularly in domains such as healthcare, where data privacy concerns and regulatory requirements often necessitate decentralized approaches. Lastly, NTK-DFL may serve as a useful paradigm for transfer learning applications in scenarios where a single, centralized source of both compute and data is not available.
(a) FEMNIST
(b) $\alpha=0.05$ MNIST
(c) $\alpha=0.1$ MNIST
Figure 8: Convergence of various methods on (a) FEMNIST, (b) Non-IID MNIST ($\alpha=0.05$), and (c) Non-IID MNIST ($\alpha=0.1$). NTK-DFL consistently outperforms all baselines.
Figure 9: The effect of static vs. dynamic topology on NTK-DFL. Solid lines correspond to a dynamic topology, whereas dotted lines correspond to a static topology. Both methods benefit from the dynamic topology and NTK-DFL outperforms DFedAvg under both topologies. Other baselines are not drawn but perform similarly to DFedAvg.
Figure 10: The effect of different vs. identical weight initialization. Solid lines correspond to the same weight initialization for all clients, whereas dotted lines correspond to different initialization. The convergence of NTK-DFL is affected less than that of DFedAvg. Other baselines are not drawn but perform similarly to DFedAvg.
@@ Appendix A NTK-DFL Algorithms
1:  Initialize weights ${\bm{w}}_{i}^{(0)}$ for each client $C_{i}$.
@@
3:     Initialize graph structure $G^{(k)}=(\mathcal{C},E^{(k)})$, specifying the neighbors $\mathcal{N}_{i}^{(k)}$ for each client $C_{i}$.
@@
0:  For each client $C_{i}$, a set of neighbors $\mathcal{N}_{i}^{(k)}$ and initial weights ${\bm{w}}_{i}^{(k)}$
1:  for each client $C_{i}\in\mathcal{C}$ in parallel do
2:     Send ${\bm{w}}_{i}^{(k)}$ to all neighbors $C_{j}\in\mathcal{N}_{i}^{(k)}$
3:     Receive ${\bm{w}}_{j}^{(k)}$ from all neighbors $C_{j}\in\mathcal{N}_{i}^{(k)}$
4:     $\bar{{\bm{w}}}_{i}^{(k)}\leftarrow\frac{1}{|\mathcal{N}_{i}^{(k)}|+1}({\bm{w}}_{i}^{(k)}+\sum_{j\in\mathcal{N}_{i}^{(k)}}{\bm{w}}_{j}^{(k)})$
5:     Send aggregated weight $\bar{{\bm{w}}}_{i}^{(k)}$ back to all neighbors $C_{j}\in\mathcal{N}_{i}^{(k)}$
@@
0:  Each client $C_{i}$ knows its neighbors $\mathcal{N}_{i}^{(k)}$ and has access to local data $\mathbf{X}_{i}$ and the aggregated weights $\bar{{\bm{w}}}_{j}^{(k)}$ for each neighbor $C_{j}\in\mathcal{N}_{i}^{(k)}$.
1:  for each client $C_{i}\in\mathcal{C}$ in parallel do
2:     Compute the Jacobian ${\bm{J}}^{(k)}_{i,i}\equiv\nabla_{\bar{{\bm{w}}}_{i}}{\bm{f}}(\mathbf{X}_{i};\bar{{\bm{w}}}_{i}^{(k)})$ using the client’s own aggregated weight $\bar{{\bm{w}}}_{i}^{(k)}$ and local data $\mathbf{X}_{i}$.
3:     for each neighbor $C_{j}\in\mathcal{N}_{i}^{(k)}$ do
4:        Compute the Jacobian ${\bm{J}}^{(k)}_{i,j}\equiv\nabla_{\bar{{\bm{w}}}_{j}}{\bm{f}}(\mathbf{X}_{i};\bar{{\bm{w}}}_{j}^{(k)})$ using the neighbor’s aggregated weight $\bar{{\bm{w}}}_{j}^{(k)}$ and client’s local data $\mathbf{X}_{i}$.
5:        Send ${\bm{J}}^{(k)}_{i,j}$, true label $\mathbf{Y}_{i}$, and function evaluation ${\bm{f}}(\mathbf{X}_{i};\bar{{\bm{w}}}_{j}^{(k)})$ to neighbor $C_{j}$.
@@
0:  Each client $C_{i}$ has access to local data $\mathbf{X}_{i}$, initial weights $\bar{{\bm{w}}}_{i}^{(k)}$, and knows its neighbors $\mathcal{N}_{i}^{(k)}$
1:  for each client $C_{i}\in\mathcal{C}$ after intra-client communication do
2:     Compute local Jacobian tensor ${\bm{J}}_{i,i}^{(k)}$ and receive ${\bm{J}}_{j,i}^{(k)}$ from each neighbor $C_{j}$
3:     Construct tensor $\bm{\mathcal{J}}_{i}^{(k)}$ from $\{{\bm{J}}_{ii}^{(k)}\}\cup\{{\bm{J}}_{ji}^{(k)}\mid C_{j}\in\mathcal{N}_{i}^{(k)}\}$
@@
6:        $\mathbf{H}^{(k)}_{i,mn}\leftarrow\frac{1}{d_{2}}\langle\bm{\mathcal{J}}_{i}^{(k)}(x_{m}),\bm{\mathcal{J}}_{i}^{(k)}(x_{n})\rangle_{F}$
@@
9:        ${\bm{f}}^{(k,t)}(\bm{\mathcal{X}}_{i})\leftarrow(\mathbf{I}-e^{\frac{\eta t}{\tilde{N_{i}}}\mathbf{H}_{i}^{(k)}})\bm{\mathcal{Y}}_{i}^{(k)}+e^{\frac{\eta t}{\tilde{N_{i}}}\mathbf{H}_{i}^{(k)}}{\bm{f}}^{(k)}(\bm{\mathcal{X}}_{i})$
10:        ${\bm{w}}_{i}^{(k,t)}\leftarrow\sum_{j=1}^{d_{2}}(\bm{\mathcal{J}}^{(k)}_{i,:j:})^{T}{\bm{R}}_{i,:j}^{(k,t)}+\bar{{\bm{w}}}_{i}^{(k)}$
@@
12:     Select ${\bm{w}}_{i}^{(k+1,0)}\leftarrow{\bm{w}}_{i}^{(k,t)}$ with the lowest loss given the residual $({\bm{f}}^{(k,t)}(\bm{\mathcal{X}}_{i})-\bm{\mathcal{Y}}_{i})$
@@
Appendix B Additional Experimental Details
B.1 Baselines
@@
NTK-FL is the only centralized baseline that we compare with. We choose a part rate which ensures that the busiest node in the centralized setting is no busier than the busiest decentralized setting. By busier, we mean the degree of the node or number of clients communicating with it. We note that NTK-FL is not an upper bound in this case due to the comparison being founded on node busyness. Evaluating NTK-FL in the same setting as the table in Figure 2, NTK-FL converges to threshold accuracy in 73, 85, and 180 communication rounds for heterogeneity settings IID, $\alpha=0.5$, and $\alpha=0.1$ respectively. D-PSGD is one of the first decentralized, parallel algorithms for distributed machine learning that allows nodes to only communicate with neighbors. DFedAvg adapts FedAvg to the decentralized setting, and DFedAvgM makes the use of SGD-based momentum and extends DFedAvg. Both use multiple local epochs between communication rounds, like vanilla FedAvg. DisPFL is a personalized federated learning approach that aims to train a global model and personalize it to each client with a local mask. In order to make the comparision fair, we report the accuracy of the global model on our test set.
B.2 Hyperparameters
@@
We perform a hyperparameter search over each baseline and select the hyperparameters corresponding to the best test accuracy. We use the $\alpha=0.1$ Fashion-MNIST test accuracy at communication round 30 as the metric for selection. This is done because the majority of comparisons take place on Fashion-MNIST in the non-IID setting. For D-PSGD, we use a learning rate of $0.1$, and a batch size of $10$ (local epochs are defined to be one in this approach). For DFedAvg, we use a learning rate of $0.1$, a batch size of $25$, and $20$ local epochs. For DFedAvgM, we use a learning rate of $0.01$, a batch size of $50$, $20$ local epochs, and a momentum of $0.9$. For DisPFL, we use a learning rate of $0.1$, a batch size of $10$, and $10$ local epochs. Following (Dai et al., 2022), we use the sparsity rate of $0.5$ for DisPFL. As for the NTK-DFL, we use a learning rate of 0.01 and search over values $t\in\{100,200,\ldots,800\}$ during the weight evolution process.
B.3 Inter-Model Variance
We describe the inter-model variance between NTK-DFL clients in Figure 7. Here, the variance is calculated as follows
$V=\frac{1}{d_{2}}\sum_{j=1}^{d_{2}}\sqrt{\sum_{i=1}^{M}[\bar{{\bm{w}}}-{\bm{w}}_{i}]_{j}^{2}}$ (6)
where we investigate the average per-parameter variance to normalize for the scale of each parameter among clients.
B.4 Further Experimental Results
@@
Figure 11: Convergence of NTK-DFL across different dynamic topologies, trained on Fashion-MNIST. NTK-DFL is evaluated with (blue) a $\kappa=5$ regular graph, (yellow) an Erdos-Renyi random graph with five mean neighbors, and (green) a ring topology, where each client is connected to two neighbors. We observe that NTK-DFL demonstrates steady convergence across different topology classes.

### v2

@@ 6 Conclusion and Future Work
In this paper, we have introduced NTK-DFL, a novel approach to decentralized federated learning that leverages the neural tangent kernel to address the challenges of statistical heterogeneity in decentralized learning settings. Our work extends NTK-based training to the decentralized setting, while discovering a unique synergy between NTK evolution and decentralized model averaging that improves final model accuracy. Our method combines the expressiveness of NTK-based weight evolution with a decentralized architecture, allowing for efficient, collaborative learning without a central server. We reduce the number of communication rounds needed for convergence, which may prove advantageous for high-latency settings or those with heavy encoding/decoding costs.
There are promising unexplored directions for NTK-DFL. For instance, extending the algorithm to training models such as CNNs, ResNets (He et al., 2016), and transformers (Vaswani et al., 2017), possibly making use of new NTK methods suited for modern architectures (Arora et al., 2019; Tirer et al., 2022; Yang, 2019). Additionally, future research could explore the application of NTK-DFL to cross-silo federated learning scenarios, particularly in domains such as healthcare, where data privacy concerns and regulatory requirements often necessitate decentralized approaches (Guo et al., 2025). Lastly, NTK-DFL may serve as a useful paradigm for transfer learning applications in scenarios where a single, centralized source of both compute and data is not available.
Impact Statement
This paper presents work aimed at advancing the field of machine learning. Our research has several potential societal consequences, none of which we believe need to be specifically highlighted in this context.
Acknowledgments
This work was supported in part by the US National Science Foundation under grants SaTC-2340856 and ECCS-2203214, and the ECE Undergraduate Research Program and the Caldwell Fellows Program at the NC State University. The views expressed in this publication are those of the authors and do not necessarily reflect the views of the National Science Foundation.
@@ Appendix A NTK-DFL Algorithms
1:  Initialize weights ${\bm{w}}_{i}^{(0)}$ for each client $i$.
@@
3:     Initialize graph structure $G^{(k)}=(\mathcal{C},E^{(k)})$, specifying the neighbors $\mathcal{N}_{i}^{(k)}$ for each client $i$.
@@
0:  For each client $i$, a set of neighbors $\mathcal{N}_{i}^{(k)}$ and initial weights ${\bm{w}}_{i}^{(k)}$
1:  for each client $i\in\mathcal{C}$ in parallel do
2:     Send ${\bm{w}}_{i}^{(k)}$ to all neighbors ${j\in\mathcal{N}_{i}^{(k)}}$
3:     Receive ${\bm{w}}_{j}^{(k)}$ from all neighbors ${j\in\mathcal{N}_{i}^{(k)}}$
4:     $\bar{{\bm{w}}}_{i}^{(k)}\leftarrow\frac{1}{{N_{i}+\sum_{j\in\mathcal{N}^{(k)}_{i}}N_{j}}}\big{(}{N_{i}{\bm{w}}_{i}^{(k)}+\sum_{j\in\mathcal{N}_{i}^{(k)}}N_{j}{\bm{w}}_{j}^{(k)}}\big{)}$
5:     Send aggregated weight $\bar{{\bm{w}}}_{i}^{(k)}$ back to all neighbors ${j\in\mathcal{N}_{i}^{(k)}}$
@@
0:  Each client $i$ knows its neighbors $\mathcal{N}_{i}^{(k)}$ and has access to local data $\mathbf{X}_{i}$ and the aggregated weights $\bar{{\bm{w}}}_{j}^{(k)}$ from each neighbor ${j\in\mathcal{N}_{i}^{(k)}}$.
1:  for each client $i\in\mathcal{C}$ in parallel do
2:     Compute the Jacobian ${\bm{J}}^{(k)}_{i,i}\equiv\nabla_{{{\bm{w}}}}f(\mathbf{X}_{i};\bar{{\bm{w}}}_{i}^{(k)})$ using the client’s own aggregated weight $\bar{{\bm{w}}}_{i}^{(k)}$ and local data $\mathbf{X}_{i}$.
3:     for each neighbor ${j\in\mathcal{N}_{i}^{(k)}}$ do
4:        Compute the Jacobian ${\bm{J}}^{(k)}_{i,j}\equiv\nabla_{{{\bm{w}}}}f(\mathbf{X}_{i};\bar{{\bm{w}}}_{j}^{(k)})$ using the neighbor’s aggregated weight $\bar{{\bm{w}}}_{j}^{(k)}$ and client’s local data $\mathbf{X}_{i}$.
5:        Send ${\bm{J}}^{(k)}_{i,j}$, true label $\mathbf{Y}_{i}$, and function evaluation $f(\mathbf{X}_{i};\bar{{\bm{w}}}_{j}^{(k)})$ to neighbor $j$.
@@
0:  Each client $i$ has access to local data $\mathbf{X}_{i}$ and initial weights $\bar{{\bm{w}}}_{i}^{(k)}$, and knows its neighbors $\mathcal{N}_{i}^{(k)}$
1:  for each client $i\in\mathcal{C}$ after intra-client communication do
2:     Compute local Jacobian tensor ${\bm{J}}_{i,i}^{(k)}$ and receive ${\bm{J}}_{j,i}^{(k)}$ from each neighbor $j$
3:     Construct tensor $\bm{\mathcal{J}}_{i}^{(k)}$ from $\{{\bm{J}}_{ii}^{(k)}\}\cup\{{\bm{J}}_{ji}^{(k)}\mid{j\in\mathcal{N}_{i}^{(k)}}\}$
@@
6:        $[\mathbf{H}^{(k)}_{i}]_{m,n}\leftarrow\frac{1}{d_{2}}\langle\bm{\mathcal{J}}_{i}^{(k)}(x_{m}),\bm{\mathcal{J}}_{i}^{(k)}(x_{n})\rangle_{F}$
@@
9:        ${\bm{f}}^{(k,t)}(\bm{\mathcal{X}}_{i})\leftarrow(\mathbf{I}-e^{-\frac{\eta t}{\tilde{N_{i}}}\mathbf{H}_{i}^{(k)}})\bm{\mathcal{Y}}_{i}^{(k)}+e^{-\frac{\eta t}{\tilde{N_{i}}}\mathbf{H}_{i}^{(k)}}{\bm{f}}^{(k)}(\bm{\mathcal{X}}_{i})$
10:        $\bar{{\bm{w}}}_{i}^{(k,t)}\leftarrow\sum_{j=1}^{d_{2}}(\bm{\mathcal{J}}^{(k)}_{i,:j:})^{T}{\bm{R}}_{i,:j}^{(k,t)}+\bar{{\bm{w}}}_{i}^{(k)}$
@@
12:     Select ${\bm{w}}_{i}^{(k+1,0)}\leftarrow\bar{{\bm{w}}}_{i}^{(k,t)}$ with the lowest loss given the residual ${\bm{f}}^{(k,t)}(\bm{\mathcal{X}}_{i})-\bm{\mathcal{Y}}_{i}$
@@
Appendix B Additional Details on Weight Evolution
In implementation, computing the matrix exponential $e^{-\frac{\eta t}{\tilde{N_{i}}}\mathbf{H}_{i}^{(k)}}$ in (4) to evolve weights can be computationally expensive. In practice, the weights are evolved according to the more general differential equation from which (4) is derived, reliant upon the linearized model approximation ${\bm{f}}({\bm{\mathcal{X}}_{i}};\bar{{\bm{w}}}_{j}^{(k,t)})\approx{\bm{f}}(\bm{\mathcal{X}}_{i};\bar{{\bm{w}}}_{j}^{(k,0)})+\nabla_{{{\bm{w}}}}{\bm{f}}(\bm{\mathcal{X}}_{i};\bar{{\bm{w}}}_{j}^{(k,0)})^{\top}(\bar{{\bm{w}}}_{j}^{(k,t)}-\bar{{\bm{w}}}_{j}^{(k,0)})$. The differential equation is as follows
$\frac{d}{dt}{\bm{f}}(\bm{\mathcal{X}}_{i};\bar{{\bm{w}}}_{j}^{(k,t)})=-{\eta}\mathbf{H}_{j}^{(k)}\nabla_{{\bm{f}}}\mathcal{L}.$ (11)
Here, $\mathcal{L}$ is the loss function. For example, for a half mean-squared error (MSE) loss, term on the right becomes the residual matrix $\nabla_{{\bm{f}}}\mathcal{L}={\bm{f}}(\bm{\mathcal{X}}_{i};\bar{{\bm{w}}}_{j}^{(k,t)})-\bm{\mathcal{Y}}_{i}$. During weight evolution, a client $j$ evolves their neighboring function evaluation from the initial condition ${\bm{f}}(\bm{\mathcal{X}}_{i};\bar{{\bm{w}}}_{j}^{(k,0)})$ to the time-evolved ${\bm{f}}(\bm{\mathcal{X}}_{i};\bar{{\bm{w}}}_{j}^{(k,t)})$ using a differential equation solver and the differential equation above. To implement (5a), we use a process similar to Yue et al. (2022) where the initial client residual is evolved over a series of timesteps specified by the user. For user-specified timesteps, the loss at that time is found using the evolved residual. Then, the best-performing weights are evolved using the left side of (5a) and selected for the next communication round.
Appendix C Additional Experimental Details
C.1 Baselines
@@
NTK-FL is the only centralized baseline that we compare with. We choose a per-round client fraction that ensures that the busiest node in the centralized setting is no busier than the busiest decentralized setting. By busier, we mean the degree of the node or the number of clients communicating with it. We note that NTK-FL is not an upper bound in this case due to the comparison being founded on node busyness, which disadvantages a centralized approach where all communication happens through a single, centralized node. Evaluating NTK-FL in the same setting as the table in Figure 2, NTK-FL converges to threshold accuracy in 73, 85, and 180 communication rounds for heterogeneity settings IID, $\alpha=0.5$, and $\alpha=0.1$, respectively. D-PSGD (Lian et al., 2017) is one of the first decentralized, parallel algorithms for distributed machine learning that allows nodes to only communicate with neighbors. DFedAvg (Sun et al., 2021) adapts FedAvg to the decentralized setting, and DFedAvgM makes the use of SGD-based momentum and extends DFedAvg. Both use multiple local epochs between communication rounds, like vanilla FedAvg. DFedSAM (Shi et al., 2023) incorporates the SAM algorithm (Foret et al., 2021) into the DFL process. DisPFL (Dai et al., 2022) is a personalized federated learning approach that aims to train a global model and personalize it to each client with a local mask. In order to make the comparision fair, we report the accuracy of the global model on our test set.
C.2 Hyperparameters
@@
We perform a hyperparameter search over each baseline and select the hyperparameters corresponding to the best validation accuracy. We use the $\alpha=0.1$ Fashion-MNIST test accuracy at communication round 30 as the metric for selection. This is done because the majority of comparisons take place on Fashion-MNIST in the non-IID setting. For D-PSGD, we use a learning rate of $0.1$, and a batch size of $10$ (local epoch count is defined to be one in this approach). For DFedAvg, we use a learning rate of $0.1$, a batch size of $25$, and $20$ local epochs. For DFedAvgM, we use a learning rate of $0.01$, a batch size of $50$, $20$ local epochs, and a momentum of $0.9$. For DisPFL, we use a learning rate of $0.1$, a batch size of $10$, and $10$ local epochs. Following Dai et al. (2022), we use the sparsity rate of $0.5$ for DisPFL. For DFedSAM, we began with the parameters suggested in Shi et al. (2023). After a hyperparameter search, we found that a radius $\rho=0.01$, $\eta=0.01$, momentum of $0.99$, learning rate decay of $0.95$, weight decay of $5\times 10^{-4}$, 5 local epochs, and a batch size of 32 yielded the best performance. Note that we used a single gossip step per round for all approaches in order to maintain a fair comparison. As for the NTK-DFL, we use a learning rate of 0.01 and search over values $t\in\{100,200,\ldots,800\}$ during the weight evolution process. For the baseline compression methods, we experimented with quantization as well as top-$k$ sparsification (see Appendix E). We found that the baseline methods, which were gradient-based, cannot withstand the aggressive sparsification applied to NTK-DFL. Therefore, we used only quantization [such as in Sun et al. (2021)] and selected the most effective value to facilitate a fair comparison.
C.3 Inter-Client Model Deviation
We plot the relationship between the accuracy and the inter-client model deviation among NTK-DFL clients in Figure 5, where the inter-client model deviation is defined for a set of weight vectors ${\bm{w}}_{1},\ldots,{\bm{w}}_{M}\in\mathbb{R}^{d}$ as follows:
$V=\frac{1}{d}\sum_{j=1}^{d}\sqrt{\sum_{i=1}^{M}(w_{i,j}-\bar{w}_{j})^{2}},\quad\bar{{\bm{w}}}=\frac{1}{M}\sum_{i=1}^{M}{\bm{w}}_{i},$ (12)
where it captures per-parameter variation in an averaged sense.
C.4 Additional Experimental Results
@@
Figure 8: Convergence of NTK-DFL across different dynamic topologies, trained on Fashion-MNIST. NTK-DFL is evaluated with a clustered graph (in yellow) with 5 neighbors per client, a $\kappa=5$ regular graph (in blue), an Erdos-Renyi random graph with five mean neighbors (in green), a ring topology (in red), and a line topology (in purple). We observe that NTK-DFL demonstrates steady convergence across different topology classes. For the ring and line topologies, convergence is a bit slower due to a sparser graph of 2 rather than 5 neighbors per client.